### v1

@@ Abstract
Robot navigation has transitioned from prioritizing obstacle avoidance to adopting socially aware navigation strategies that accommodate human presence. As a result, the recognition of socially aware navigation within dynamic human-centric environments has gained prominence in the field of robotics. Although reinforcement learning technique has fostered the advancement of socially aware navigation, defining appropriate reward functions, especially in congested environments, has posed a significant challenge. These rewards, crucial in guiding robot actions, demand intricate human-crafted design due to their complex nature and inability to be automatically set. The multitude of manually designed rewards poses issues with hyperparameter redundancy, imbalance, and inadequate representation of unique object characteristics. To address these challenges, we introduce a transformable gaussian reward function (TGRF). The TGRF significantly reduces the burden of hyperparameter tuning, displays adaptability across various reward functions, and demonstrates accelerated learning rates, particularly excelling in crowded environments utilizing deep reinforcement learning (DRL). We introduce and validate TGRF through sections highlighting its conceptual background, characteristics, experiments, and real-world application, paving the way for a more effective and adaptable approach in robotics. The complete source code is available on https://github.com/JinnnK/TGRF
@@ 1 Introduction
Over the years, the robotics field has shown a persistent interest in robot navigation. Initially, research efforts focused on basic obstacle avoidance and random navigation strategies [1-3]. Advances in navigation techniques led to simultaneous localization and mapping (SLAM) [4-6], where robots estimate their positions and create maps to move effectively. Strategies expanded to address dynamic environments [7-17], as robotics consistently pursued advancements in navigation.
@@
However, despite the integration of robots into human environments, effectively avoiding collisions with humans remains a significant challenge. While safety, performance, and navigation in static environments were previously emphasized [18], the field has now shifted its focus to socially aware navigation, recognized as crucial [12-14, 16, 17]. Socially aware navigation integrates perception, intelligence, and behavior to adhere to social norms [19, 20], necessitating the critical ability to differentiate between humans and static objects.
@@
Figure 1: The robot’s actions when the adequate reward functions are used or not. The red shapes represent penalties, and the yellow arrows indicate the robot’s actions. Penalties are imposed when the robot is in proximity to humans, within their surroundings, or moving in their direction.
@@
To deal with this challenge, two main research directions emerged: reactive navigation [9-11, 15] and navigation utilizing reinforcement learning (RL) [21-24]. Reactive navigation responds to real time sensor data, with limitations in foreseeing future movements. RL uses markov decision process (MDP) and deep reinforcement learning (DRL) [12-14, 16, 17, 25], utilizing deep neural networks for well-informed decisions and enabling robots to navigate safely in human environments.
@@
However, the challenges in defining these reward functions become particularly evident in crowded environments where robots navigate [21, 26-28]. These reward functions essentially serve as the guiding principle steering the actions of agents by evaluating the potential value of each action. As a result, human-crafted rewards become indispensable as they can’t be automatically set. However, as demonstrated in Fig. 1, inadequately designed reward functions can induce risky behaviors in human-robot interactions. Moreover, the manual design of numerous such rewards presents several critical issues.
@@
First, the proliferation of distinct reward functions for various components in crowded environments necessitates a redundant number of hyperparameters [12-14, 16, 17, 25, 29-31]. Each reward demands tailored functions that align with its specific attributes, be it the distance from humans, the direction towards the goal, or even human intentions. This surplus of reward functions results in excessive fine-tuning of hyperparameters, leading to imbalances in the rewards. These imbalances can inadvertently steer robots towards humans, increasing the risk of collisions [12]. Such disparities in action priorities, caused by reward imbalances pose a significant challenge in tuning hyperparameters and identifying inadequate rewards.
@@
Second, the fixed form of each reward function is not temporally and experimentally inefficient. The utilization of static context specific reward functions has been observed [12, 14, 16, 17, 32-35]. However, these often fall short in adequately representing the its unique characteristics. Even when the same formula is employed, the diverse attributes may not be accurately captured. This discrepancy demands extensive empirical design and experimentation with specific reward functions to achieve higher performance.
@@
Third, the hurdle can extend to effective learning [27]. Crafting appropriate reward functions with the correct number of hyperparameters remains a significant challenge, leading to collisions and hindering robot learning.
@@
This study proposes a transformable gaussian reward function (TGRF), aiming to guide robots along safer routes. This approach brings several crucial contributions. (1) The less number of hyperparameters significantly alleviates the burden of parameter tuning, expediting the search for an optimal reward function. (2) TGRF demonstrates adaptability to various reward functions through dynamic shape adjustments. Such adaptability stands in stark contrast to previous models [13], which often required extensive redesigns for shape changes. (3) TGRF shows accelerated learning rates, excelling notably in crowded environments, effectively harnessing the potential of DRL.
@@
To demonstrate the performance of TGRF, we introduce key points in reward shaping and relevant papers for comparison in experiments in Section II. In Section III, we present background knowledge, characteristics of TGRF, and introduces the reward functions using TGRF. In Section IV, we showcase two experiments conducted to demonstrate the performance of TGRF, and present the results of this study’s application in real environments in Section V. Finally, we concludes the paper in Section VI.
@@ 2 Related Works
2.1 Integration of Prior Knowledge through Human Delivered Reward Functions
Reinforcement learning is a machine learning approach operating within MDP [21], where an agent interacts with a specific environment, receiving feedback in the form of rewards. The primary objective is to achieve the maximum cumulative reward, driving the learning of policies. The reward function significantly influences the agent’s decision making process, steering it towards converging to optimal policies.
However, a prevalent issue arises as agents typically lack complete information about all aspects of the environment [36]. In vast state spaces, transitions between states and rewards might be unknown or stochastic. Moreover, the agent remains unaware of the consequences and outcomes of actions until they are executed.
As a result, in such scenarios, agents require substantial experience to converge to optimal policies for complex tasks in the absence of prior knowledge. To address this challenge, research in reinforcement learning explores reward shaping, aiming to guide agents towards making better decisions at appropriate times through suitable reward values [27, 28]. This approach aims to significantly reduce the learning time by fostering the convergence to optimal policies without explicit prior knowledge.
Previous studies have extensively explored incorporating prior knowledge into reward functions [26, 28, 37-40]. However, crafting reward functions encompassing general prior knowledge, such as apprehension concerning collision risks based on proximity to humans or progress relative to the destination, is challenging due to various environmental and psychological factors. These factors make it impossible to express such knowledge simply through mathematical formulations.
To address this challenge, recent research has focused on utilizing inverse reinforcement learning (IRL), where humans intervene at intermediate stages to provide rewards [41]. In addition, a study utilizing natural language to communicate intermediate reward functions to agents have emerged [42]. These studies involve humans evaluating the actions of robots as rewards and assigning them accordingly. As a result, they have demonstrated the transmission of rewards imbued with prior knowledge to agents during learning, thereby accelerating the learning process and enhancing algorithm performance.
However, due to their reliance on human intervention, these approaches are unsuitable for environments requiring extensive learning or complex tasks without direct human involvement. Therefore, there is a growing need for research on reward shaping that can consider prior knowledge and deliver high performance without direct human intervention, especially in environments demanding substantial learning.
@@ 2.2 Reward Function Analysis for Human Avoidance in Robot Navigation
Recent research has been actively exploring methods for robot navigation in environments where humans and robots coexist [12-14, 16, 17, 25, 29-31]. Their reward functions commonly employ different formulas based on objectives, all without direct human intervention, and can broadly be classified into four types. These include rewards categorized as follows: reaching the destination, $r_{goal}(s_{t})$ ; collision avoidance with humans, $r_{col}(s_{t})$; distance from humans, $r_{disc}(s_{t})$; and distance from the destination, $r_{pot}(s_{t})$.
@@
$r_{goal}(s_{t})$ and $r_{col}(s_{t})$ usually assume consistent values, while $r_{disc}(s_{t})$ and $r_{pot}(s_{t})$ encompass numerous formulas reflecting prior knowledge, such as linear, L2 norm, or exponential functions. For instance, $r_{dist}(s_{t})$ consistently imposes a larger penalty as the distance between humans and robots diminishes. This design aligns with the psychological theory of Proxemics [20], which evaluates discomfort based on interpersonal distances, integrating prior knowledge about the potential discomfort associated with varying distances between humans and robots.
@@
In addition, $r_{pot}(s_{t})$ incentivizes the robot’s faster arrival at the destination by applying rewards or penalties based on changes in the L2 norm distance between the robot and the destination. These approaches reflect rational strategies by integrating prior knowledge concerning discomfort levels associated with distances (Proxemics) and apprehensions regarding collision probabilities based on proximity to humans. Moreover, they encourage robots to expedite their arrival at the destination through rewards/penalties based on the changing distance from the destination using the L2 norm.
@@
However, papers related to reward shaping and reinforcement learning argue that it is crucial to verify if the rewards take appropriate forms and maintain suitable proportions [27, 28]. If the shapes of rewards are inadequate for the objectives or overly biased, the robot may steer its learning process in a direction not intended by the algorithm, potentially leading to the freezing robot problem [42]. In addition, an excessive number of hyperparameters might hinder the quest for optimal performance.
@@
The aforementioned prior studies experimentally determined reward functions and the count of hyperparameters. As a result, some were excessively simplistic, preventing researchers from intuitively adjusting rewards through hyperparameters, while others exhibited complex structures that hindered straightforward modifications of hyperparameters. This resulted in significant time consumption to achieve optimal performance and limitations in adjusting inadequate rewards, necessitating redesigns of the reward function.
@@
For instance, in [13], researchers designed simple reward functions with redundant hyperparameters. This led to a substantial nine fold gap between $r_{pred}(s_{t})$ and $r_{disc}(s_{t})$. This results in situations in which the robot favors actions with smaller penalties from $r_{disc}(s_{t})$ over larger penalties from $r_{pred}(s_{t})$, thereby resulting in intrusions and collisions.
@@
This directly impacts the learning process, rendering the task of identifying the appropriate reward function and hyperparameters more challenging and requiring formula modifications.
@@
Therefore, this study proposes TGRF, allowing for intuitive and versatile applications with less hyperparameters. By enabling researchers to intuitively adjust rewards, it reduces the time required for exploring suitable reward functions and ensures optimal performance by finely tuning reward balances. To substantiate this claim, we directly compare the reward function utilized in socially aware navigation (SCAN) [16], decentralized structural-RNN (DS-RNN) [17], gumbel social transformer + human-human attention (GST+HH Attn) [13, 43], and crowd aware memory based reinforcement learning (CAM-RL) [31].
@@ 3 Suggested Reward Function
In section III-A, we briefly introduce background knowledge regarding the model. Subsequently, in section III-B, we elaborate on the proposed TGRF in this paper. Lastly, in III-B-3, we describe the application of TGRF to the reward functions within the environment and model of [13].
@@ 3.1.1 Markov decision process (MDP) and navigation methods
Figure 2: Diagram of an MDP with the Navigation model applied.
@@
The robot utilizes a learning based model called GST + HH Attn [13], which employs an Attention mechanism [45] for navigation based on the past and current positions of humans to predict future positions.
In the GST + HH Attn model, the robot begins at $S_{0}$ and receives as state information: the robot’s current position ($p_{x}$,$p_{y}$), velocity ($v_{x}$,$v_{y}$), destination ($g_{x}$,$g_{y}$), maximum velocity $v_{max}$ , angle $\theta$, and robot radius $\rho$, along with the set of $i-th$ person’s current position information ($p_{x}^{i}$,$p_{y}^{i}$) at time $t$, denoted as $u_{i}^{t}$. $i$ can vary from 0 to a maximum number of people, $n_{max}$. However, due to limitations in LiDAR range, the total number of people n in the received state by the robot might differ from $n_{max}$ and can vary at each time t. In addition, the robot remembers the past positions of detected people up to a constant M steps, using the trajectory prediction algorithm (GST) to nonlinearly predict $\hat{u}_{i}^{t+1:t+K}$, the predicted positions of individuals from time t+1 to t+K.
Therefore, the GST + HH Attn model leverages the set of predicted positions $\hat{U}_{i}^{t+1:t+K}=[\hat{u}_{0}^{t+1:t+K},\hat{u}_{1}^{t+1:t+K},...,\hat{u}_{n}^{t+1:t+K}]$ up to constant K steps and the set of current positions $U^{t}=[u_{0}^{t},u_{1}^{t},u_{2}^{t},…,u_{n}^{t}]$ of people to predict the overall dynamics of people through an Attention mechanism. It employs a type of recurrent neural network (RNN) called GRU to detect and process longterm dependencies, enabling a broader field of view [46]. The model focuses on nearby individuals while selecting actions. Furthermore, it receives feedback and updates its state for the next step based on the chosen actions. Besides, the study employs five learning based methods using the same model:
  • DS-RNN: A model utilizing an RNN. It does not predict trajectories.
  • No pred + HH Attn: Attention based model, excluding trajectory prediction ($r_{pred}=0$).
  • Const vel + HH Attn: An experimental case assumes that the trajectory prediction algorithm predicts trajectories to move at a constant velocity.
  • Truth + HH Attn: An experiment assuming that the robot predicts the actual human trajectory.
  • GST + HH Attn: Scenarios in which the robot predicts the human trajectory nonlinearly using the GST.
@@
Each human was randomly generated with a starting location and destination. Humans exchange their location information $U^{t}=[u_{0}^{t},u_{1}^{t},u_{2}^{t},…,u_{n}^{t}]$ with each another and calculate velocities based on the positions, engaging in collision avoidance by altering their speed and direction using reactive based methods such as ORCA and SF [9-11].
Similar to robots, humans were designated destinations, and they aim to reach or modify them with a predetermined probability. Humans also possess attributes such as a size, speed, and maximum speed.
However, humans were assumed to disregard a robot’s location for two primary reasons. First, in crowded environments, human reactions to robots lead to subtle movements resulting in negligible positional changes. Even, rapid shifts are infeasible in densely populated settings. Second, encapsulating human dynamics within a mathematical model is challenging and can impede the learning process of robots. The diversity of the manners in which humans react to robots, along with associated uncertainties, makes modeling impractical. Furthermore, even if such modeling were feasible, it would interfere with the robot’s learning. As a result, humans do not perceive the robot’s location and ignore it.
Therefore, humans make their next actions based on their own characteristics, and information about the current positions and velocities of others.

### v2

@@ Abstract
Robot navigation has transitioned from prioritizing obstacle avoidance to adopting socially-aware navigation strategies that accommodate human presence. Consequently, socially-aware navigation in dynamic human-centric environments has gained prominence in the field of robotics. Although reinforcement learning techniques have fostered the advancement of socially-aware navigation, defining appropriate reward functions, particularly in congested environments, poses a significant challenge. These rewards, crucial for guiding robot actions, demand intricate human-crafted design owing to their complex nature and inability to be set automatically. The multitude of manually designed rewards poses issues such as hyperparameter redundancy, imbalance, and inadequate representation of unique object characteristics. To address these challenges, we introduce a transformable Gaussian reward function (TGRF). The TGRF significantly reduces the burden of hyperparameter tuning, displays adaptability across various reward functions, and demonstrates accelerated learning rates, particularly in crowded environments utilizing deep reinforcement learning (DRL). We introduce and validate the TGRF in sections highlighting its conceptual background, characteristics, experiments, and real-world applications, paving the way for a more effective and adaptable approach to robotics. The complete source code is available at https://github.com/JinnnK/TGRF.
@@ 1 Introduction
Over the years, robotics has shown a persistent interest in robot navigation. Initially, research focused on basic obstacle avoidance and random navigation strategies [1-3]. Advances in navigation techniques have led to simultaneous localization and mapping (SLAM) [4-6], wherein robots estimate their positions and create maps for effective movement. Strategies have expanded to address dynamic environments [7-17], as robotics has consistently pursued advancements in navigation.
@@
However, despite the integration of robots into human environments, the effective avoidance of collisions remains a significant challenge. While safety, performance, and navigation in static environments have been emphasized [18], the field has shifted its focus to socially-aware navigation, which is recognized as crucial [12-14, 16, 17]. Socially-aware navigation integrates perception, intelligence, and behavior to adhere to social norms [19, 20], necessitating the critical ability to differentiate between humans and static objects.
@@
Figure 1: Robot’s actions when adequate reward functions are used or not. Red shapes represent penalties, and yellow arrows indicate robot’s actions. Penalties are imposed when the robot is in proximity to humans, within their surroundings, or moving in their direction.
@@
To address this challenge, two main research directions have emerged: reactive navigation [9-11, 15] and navigation utilizing reinforcement learning (RL) [21-24]. Reactive navigation responds to real-time sensor data, with limitations in predicting future movements. RL employs the Markov decision process (MDP) and deep reinforcement learning (DRL) [12-14, 16, 17, 25] to leverage deep neural networks for well-informed decisions and enable robots to navigate safely in human environments.
@@
However, the challenges in defining these reward functions become particularly evident in crowded environments where robots navigate [21, 26-28]. These reward functions essentially serve as the guiding principles for steering the actions of agents by evaluating the potential value of each action. As a result, human-crafted rewards have become indispensable because they cannot be set automatically. However, as demonstrated in Fig. 1, inadequately designed reward functions can induce risky behaviors in human–robot interactions. Moreover, the manual design of numerous rewards presents several critical issues.
@@
First, the proliferation of distinct reward functions for various components in crowded environments necessitates a redundant number of hyperparameters [12-14, 16, 17, 25, 29-31]. Each reward demands tailored functions that align with its specific attributes such as distance from humans, direction toward the goal, or even human intentions. This surplus of reward functions results in excessive fine-tuning of the hyperparameters, leading to imbalances in the rewards. These imbalances can inadvertently steer robots toward humans, thereby increasing the risk of collisions [12]. Such disparities in action priorities caused by reward imbalances pose a significant challenge in tuning hyperparameters and identifying inadequate rewards.
@@
Second, the fixed form of each reward function is neither temporally nor experimentally efficient. Static context-specific reward functions were used [12, 14, 16, 17, 32-35]. However, they often fail to adequately represent their unique characteristics. Even when the same formula is employed, diverse attributes may not be captured accurately. This discrepancy requires extensive empirical design and experimentation with specific reward functions to achieve higher performance.
@@
Third, the hurdles can be extended to effective learning [27]. Crafting appropriate reward functions with the correct number of hyperparameters remains a significant challenge, which leads to collisions and hinders robot learning.
@@
This paper proposes a transformable Gaussian reward function (TGRF) that aims to guide robots along safer routes. This approach makes several crucial contributions to the literature. (1) A smaller number of hyperparameters significantly alleviates the burden of parameter tuning, thus expediting the search for the optimal reward function. (2) The TGRF demonstrated adaptability to various reward functions through dynamic shape adjustments. Such adaptability is in stark contrast to previous models [13], which often require extensive redesigns for shape changes. (3) The TGRF exhibits accelerated learning rates, notably in crowded environments, effectively harnessing the potential of DRL.
@@
To demonstrate the performance of the TGRF, we introduce the key points in reward shaping and relevant papers for comparison in the experiments in Section II. In Section III, we present background knowledge and characteristics of the TGRF and introduce the reward functions using the TGRF. In Section IV, we present two experiments conducted to demonstrate the performance of the TGRF and present the results of this study’s application in real environments; finally, we conclude the paper in Section VI.
@@ 2 Related Works
2.1 Integration of Prior Knowledge through Human-Delivered Reward Functions
RL is a machine learning approach operating within the MDP [21], where an agent interacts with a specific environment and receives feedback in the form of rewards. The primary objective is to achieve the maximum cumulative reward that drives policy learning. The reward function significantly influences the agent’s decision-making process, steering it toward convergence to optimal policies.
However, a prevalent issue arises because agents typically lack complete information regarding all aspects of the environment [36]. In vast state spaces, the transitions between states and rewards may be unknown or stochastic. Moreover, the agent remains unaware of the consequences and outcomes of the actions until they are executed.
Consequently, in such scenarios, agents require substantial experience to converge to optimal policies for complex tasks in the absence of prior knowledge. To address this challenge, research on RL has explored reward shaping, aiming to guide agents toward making better decisions at appropriate times using suitable reward values [27, 28]. This approach aims to significantly reduce the learning time by fostering convergence to optimal policies without explicit prior knowledge.
Previous studies extensively explored incorporating prior knowledge into reward functions [26, 28, 37-40]. However, crafting reward functions encompassing general prior knowledge, such as the apprehension of collision risks based on proximity to humans or progress relative to the destination, is challenging because of various environmental and psychological factors. These factors render it impossible to express knowledge simply through mathematical formulations.
To address this challenge, recent research has focused on utilizing inverse reinforcement learning (IRL), wherein humans intervene at the intermediate stages to provide rewards [41]. In addition, a study utilizing natural language to communicate intermediate reward functions with agents has emerged [42]. These studies involved humans evaluating the actions of robots as rewards and assigning them accordingly. They demonstrated the transmission of rewards imbued with prior knowledge to agents during learning, thereby accelerating the learning process and enhancing algorithm performance.
However, because of their reliance on human intervention, these approaches are unsuitable for environments that require extensive learning or complex tasks without direct human involvement. Therefore, there is a growing need for research on reward shaping that considers prior knowledge and delivers high performance without direct human intervention, particularly in environments that demand substantial learning.
@@ 2.2 Reward Function Analysis for Human Avoidance in Robot Navigation
Recent studies have actively explored methods for robot navigation in environments where humans and robots coexist [12-14, 16, 17, 25, 29-31]. Their reward functions commonly employ different formulas based on objectives without direct human intervention and can be broadly classified into four types. These include rewards categorized as follows: reaching the destination, $r_{goal}(s_{t})$; collision avoidance with humans, $r_{col}(s_{t})$; distance from humans, $r_{disc}(s_{t})$; and distance from the destination, $r_{pot}(s_{t})$.
@@
$r_{goal}(s_{t})$ and $r_{col}(s_{t})$ typically assume consistent values, whereas $r_{disc}(s_{t})$ and $r_{pot}(s_{t})$ encompass numerous formulas that reflect prior knowledge, such as linear, L2 norm, or exponential functions. For instance, $r_{dist}(s_{t})$ consistently imposes a larger penalty as the distance between humans and robots diminishes. This design aligns with the psychological theory of proxemics [20], which evaluates discomfort based on interpersonal distances, and integrates prior knowledge about the potential discomfort associated with varying distances between humans and robots.
@@
In addition, $r_{pot}(s_{t})$ incentivizes the robot’s faster arrival at the destination by applying rewards or penalties based on changes in the L2 norm distance between the robot and the destination. These approaches reflect rational strategies by integrating prior knowledge of discomfort levels associated with distances (proxemics) and apprehensions regarding collision probabilities based on proximity to humans. Moreover, they encourage robots to expedite their arrival at their destination through rewards/penalties based on the changing distance from the destination, using the L2 norm.
@@
However, studies related to reward shaping and RL argue that it is crucial to verify whether rewards take appropriate forms and maintain suitable proportions [27, 28]. If the shapes of the rewards are inadequate for the objectives or overly biased, the robot may steer its learning process in a direction not intended by the algorithm, potentially leading to the freezing robot problem [42]. In addition, an excessive number of hyperparameters may hinder the search for optimal performance.
@@
The aforementioned studies experimentally determined the reward functions and counts of the hyperparameters. As a result, some were excessively simplistic, preventing researchers from intuitively adjusting rewards through hyperparameters, whereas others exhibited complex structures that hindered the straightforward modification of hyperparameters. This resulted in significant time consumption to achieve optimal performance and limitations in adjusting inadequate rewards, necessitating a redesign of the reward function.
@@
For instance, in [13], researchers designed simple reward functions with redundant hyperparameters. This resulted in a substantial nine-fold difference between $r_{pred}(s_{t})$ and $r_{disc}(s_{t})$. This leads to situations in which the robot favors actions with smaller penalties from $r_{disc}(s_{t})$ over larger penalties from $r_{pred}(s_{t})$, thereby resulting in intrusion and collision.
@@
This directly affects the learning process, rendering the task of identifying the appropriate reward function and hyperparameters more challenging, and requiring formula modification.
@@
Therefore, this paper proposes the TGRF, which allows for intuitive and versatile applications with fewer hyperparameters. Enabling researchers to adjust rewards intuitively reduces the time required to explore suitable reward functions and ensures optimal performance by finely tuning reward balances. To substantiate this claim, we directly compared the reward function used in socially-aware navigation (SCAN) [16], decentralized structural-recurrent neural network (RNN) (DS-RNN) [17], Gumbel social transformer + human–human attention (GST+HH Attn) [13, 43], and crowd-aware memory-based RL (CAM-RL) [31].
@@ 3 Suggested Reward Function
In Section III-A, we briefly introduce background knowledge regarding the model. In Section III-B, we elaborate on the proposed TGRF. Finally, in Section III-B-3, we describe the application of the TGRF to the reward functions within the environment and model of [13].
@@ 3.1.1 Markov decision process (MDP) and navigation methods
Figure 2: Diagram of an MDP with navigation model applied
@@
The robot utilizes a learning-based model called GST + HH Attn [13], which employs an attention mechanism [45] for navigation based on the past and current positions of humans to predict future positions.
In the GST + HH Attn model, the robot begins at $S_{0}$ and receives as state information: the robot’s current position ($p_{x}$,$p_{y}$), velocity ($v_{x}$,$v_{y}$), destination ($g_{x}$,$g_{y}$), maximum velocity $v_{max}$ , angle $\theta$, and robot radius $\rho$, along with the set of $i-th$ person’s current position information ($p_{x}^{i}$,$p_{y}^{i}$) at time $t$, denoted as $u_{i}^{t}$. $i$ can vary from zero to the maximum number of people, $n_{max}$. However, owing to the limitations in the light detection and ranging (LiDAR) range, the total number of people n in the state received by the robot might differ from $n_{max}$ and can vary at each time t. In addition, the robot remembers the past positions of the detected people up to a constant M step using the trajectory-prediction algorithm (GST) to nonlinearly predict $\hat{u}_{i}^{t+1:t+K}$, the predicted positions of individuals from time t+1 to t+K.
Therefore, the GST + HH Attn model leverages the set of predicted positions $\hat{U}_{i}^{t+1:t+K}=[\hat{u}_{0}^{t+1:t+K},\hat{u}_{1}^{t+1:t+K},...,\hat{u}_{n}^{t+1:t+K}]$ up to constant K steps and the set of current positions $U^{t}=[u_{0}^{t},u_{1}^{t},u_{2}^{t},…,u_{n}^{t}]$ of people to predict their overall dynamics through an attention mechanism. It employs a type of RNN called a GRU to detect and process long-term dependencies, thus enabling a broader field of view [46]. This model focuses on nearby individuals when selecting their actions. Furthermore, it receives feedback and updates its state for the next step based on the selected actions. In addition, this study employed five learning-based methods using the same model:
  • DS-RNN: A model utilizing an RNN. However, it did not predict the trajectories.
  • No pred + HH Attn: Attention-based model excluding trajectory prediction ($r_{pred}=0$).
  • Const vel + HH Attn: The experimental case assumes that the trajectory-prediction algorithm predicts the trajectories to move at a constant velocity.
  • Truth + HH Attn: This experiment assumes that the robot predicts the actual human trajectory.
  • GST + HH Attn: Scenarios in which the robot predicts the human trajectory nonlinearly using GST.
@@
Each human was randomly assigned a starting location and destination. Humans exchange their location information $U^{t}=[u_{0}^{t},u_{1}^{t},u_{2}^{t},…,u_{n}^{t}]$ with each other and calculate velocities based on their positions, engaging in collision avoidance by altering their speed and direction using reactive-based methods such as ORCA and SF [9-11].
Similar to robots, humans are designated destinations that aim to reach or modify using a predetermined probability. Humans possess attributes such as size, speed, and maximum speed.
However, the humans were assumed to disregard the robot’s location for two primary reasons: First, in crowded environments, human reactions to robots lead to subtle movements resulting in negligible positional changes. Even rapid shifts are infeasible in densely populated settings. Second, encapsulating human dynamics within a mathematical model is challenging and can impede the learning process of robots. The diversity of the manners in which humans react to robots, along with the associated uncertainties, makes modeling impractical. Furthermore, even if such modeling were feasible, it would interfere with the robot’s learning. Consequently, humans either do not perceive the robot’s location or ignore it.
Therefore, humans perform subsequent actions based on their own characteristics and information about the cur-rent positions and velocities of others.